OpenStudy (anonymous):

Simplify 3t^3 * 4t * 2t^4 A) 24t^12 B) 24t^7 C) 24t^8 D) 24t^13 I used * for times :) please explain to me how to do this also :)

OpenStudy (anonymous):

when you multiply variables with exponents you add the exponents together

OpenStudy (anonymous):

but only if the variables are the same

OpenStudy (anonymous):

So it would be B?

OpenStudy (anonymous):

well if you add up all the exponents what do you get?

OpenStudy (anonymous):

keep in mind that t is the same thing as t^1

OpenStudy (anonymous):

I got 10^4 but I know that wrong :/

OpenStudy (anonymous):

now I got 24^7

OpenStudy (anonymous):

ok well \[3t ^{3}*4t=12t ^{3+1}=12t ^{4}\]

OpenStudy (anonymous):

do you follow so far?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok well 4*3 is just 12 right? then we just add the exponents. t^3+1=t^4

OpenStudy (anonymous):

so we get 12t^4

OpenStudy (anonymous):

ok I get that part

OpenStudy (anonymous):

now I have 24t^8 am I right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

WooooH!!!! thank you!!!!! :)